OpenStudy (anonymous):

(1+tanx)(1-tanx)+sec^2x

OpenStudy (anonymous):

the answer is two, i am completely lost.

OpenStudy (ankit042):

1-tan^x +sec^2x Now can you solve?

OpenStudy (ankit042):

HINT use trigonometric identity

OpenStudy (anonymous):

ankit, so we have tangent raised to the x power?

OpenStudy (ankit042):

sorry made a typo it is tan^2x

OpenStudy (mathstudent55):

1-tan^2 x +sec^2 x = 2

OpenStudy (anonymous):

\[1-\tan ^2x\]

OpenStudy (anonymous):

?

OpenStudy (ankit042):

yes

OpenStudy (anonymous):

should i do sin^2x over cosine ^2x

OpenStudy (ankit042):

NO there is one for tanx

OpenStudy (ankit042):

1+tan^2x = sec^2x

OpenStudy (anonymous):

so we bring the sec to the otherside? Would we attach a negative to it?

OpenStudy (ankit042):

Try replacing sec^2 with 1+tan^2

OpenStudy (anonymous):

oh, lol i see what you are saying

OpenStudy (ankit042):

:)

OpenStudy (anonymous):

so im left with two? the tangents cancel out?

OpenStudy (anonymous):

wow, i over looked this problem if i am correct. Thank you for the help ankit. :)

OpenStudy (ankit042):

no problem!

OpenStudy (jhannybean):

\[\large (1+\tan(x))(1-\tan(x))+\sec^2(x)\]\[\large 1-\tan^2(x) +\color{green}{\sec^2(x)}\]\[\large \color{green}{1+\cancel{\tan^2(x)}}+1-\cancel{\tan^2{x}}\]

OpenStudy (jhannybean):

Omg the lag. Good job Ankit.

OpenStudy (mathstudent55):

\( (1+\tan x)(1-\tan x)+sec^2x = 2 \) \( 1 - \tan^2 x + \sec^2 x = 2 \) Use the identity : \( \tan^2 x + 1 = \sec^2 x \) to replace \( \sec^2 x \) with \( \tan^2 x + 1 \) \( 1 - \tan^2 x + \tan^2 x + 1= 2 \) \(2 = 2\)